OpenStudy (anonymous):

\frac { \sqrt [ 3 ]{ x } -\sqrt [ 3 ]{ 2 } }{ x-2 }

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\frac { \sqrt [ 3 ]{ x } -\sqrt [ 3 ]{ 2 } }{ x-2 } \]

OpenStudy (anonymous):

please help.

OpenStudy (anonymous):

the answer is supposed to be \[\frac { 1 }{ 3 } \sqrt [ 3 ]{ 4 } \]

OpenStudy (anonymous):

Plugging in 2 you get an indeterminant form, so you should be able to use l'hopitals rule

OpenStudy (turingtest):

hint\[(a^3-b^3)=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

haven't gotten to l'hopitals rule yet

OpenStudy (turingtest):

so use my hint :)

OpenStudy (anonymous):

yeah, thank you :))

OpenStudy (anonymous):

finished. :P

OpenStudy (turingtest):

:)